OpenStudy (anonymous):

Gavin and Seiji both worked hard over the summer. Together, they earned a total of $425. Gavin earned $25 more than Seiji. How much did each of them earn? (a) Write a system of two equations with two variables to model this problem. (b) Use substitution or the elimination method to solve the system. (c) Graph both equations. (d) Answer the question. I already did all except C. I need help for C! Can somebody explain this to me please? (lol it rhymes.) And heres my work. G=Gavin, S= Seiji Gavin=Seiji+25;g+s=425 g+s=425 s+25+s=425 2s+25=425 2s+25+−25=425+−25 2s=400 2s/2=400/2 s=200 g=225

OpenStudy (anonymous):

@satellite73 sorryyy, i need help pls

OpenStudy (anonymous):

how do i graph both equations?

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

So, you need help in graphing these two equations? g=s+25 g+s=425

OpenStudy (anonymous):

correct! but 25 is 225.

OpenStudy (mathmale):

Let's look at the first equation: g=s+225. Suppose you let s = 0. What would g be? Now suppose you let g = 0. What would s be?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

NVM of the other answer. yes its correct

OpenStudy (mathmale):

No, it's 25, not 250.

OpenStudy (mathmale):

Let's look at the first equation: g=s+25. Suppose you let s = 0. What would g be? Now suppose you let g = 0. What would s be?

OpenStudy (mathmale):

When s=0, g= ? When g=0, s= ?

OpenStudy (anonymous):

s = 425? g = 25

OpenStudy (mathmale):

I wanted to look at the first equation first. Our goal is to graph this line, and to do so I am asking you to find the vertical and horizontal intercepts. Let's look at the first equation: g=s+25. Suppose you let s = 0. What would g be? Now suppose you let g = 0. What would s be?

OpenStudy (mathmale):

If you let s=0 here, we get g=0+25. If you let g = 0, we get 0=s+25. Solving for s, s=-25 when g=0.

OpenStudy (mathmale):

Solving g=0+25 for g, we get g=25. Let the points be (g,s). Then your vertical intercept is (0,25). Have you done this type of calculation before?

OpenStudy (anonymous):

i have, but i don't really remember, since i dont practice alot lol

OpenStudy (anonymous):



OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

@mathmale dont forget about me senpai ;c

OpenStudy (mathmale):

I'll respond in another minute or two.

OpenStudy (mathmale):

We need to plot the two points (0,-25) and (25,0).

OpenStudy (mathmale):

Note that I am assuming that g is first and s is second: (g,s) = (0, -25).